Suppose that after making 50 payments, the interest rate changes to J2=9% p.a.:

A. Convert the interest rate J2=9% to J12 equivalent (2 marks
B. Assuming that the family seeks to accept the change in interest rates, what would be their new payment based on the new interest rate? (5 marks)
C. Assuming that the family seeks to continue their initial monthly payment calculated in part I, how many full payments would be required to pay off the loan and what would be the final concluding smaller payment one period later? (9 marks
1
Expert's answer
2020-06-09T19:00:53-0400


A. "J1=(1+\frac{j2}{2})^2-1=(1+0.045)^2-1=0.092025"


"i=12\times[(1+\frac{0.092025.}{1})^{1/12}-1]"

i=0.088357 8.8357%

B.

The remainder can be found in the formula

B = (PMT/R) x (1 - (1/(1+R)^N) 

"242 720=A\times\frac{1-(1.007363)^{-310}}{0.007363}"

A=1 992.12

C.

"\frac{727056}{2019.6}=360"

360 full payments required to repay the loan


"727 056-2019*359=2019.6"